if an object placed at R_ above the earth surface then find acceleration due to gravity at that point 4

Grade:9

1 Answers

Dhruv
6 years ago
g/4
 
We know
g = (GM)/ R2   …............ (1) and   at height h,  g’ = g = (GM)/ (R+h)2   …............ (2)
 
Dividing eq(2) by eq(1), we get g’/g = (R)2/ (R+h)        put h = R     we get the answer g’/g = ¼     g’ = g/4